# Neutral-current background induced by atmospheric neutrinos at large liquid-scintillator detectors: III. Quantitative calculations for reactor neutrinos

### v1

@@ Abstract
Atmospheric neutrinos contribute significantly to irreducible backgrounds through their neutral-current (NC) interactions with C nuclei in liquid-scintillator detectors, impacting diffuse supernova neutrino background, nucleon decay, and reactor neutrinos. This paper extends our prior work [1] by systematically studying the NC backgrounds towards the MeV region of reactor neutrinos. We employ contemporary neutrino generator models from GENIE and NuWro for calculations, with a focus on predicting NC background in experimental searches for inverse-beta-decay signals below 100 MeV visible energy. We estimate the systematic uncertainty to our estimation of the NC background using various data-driven neutrino generator models, addressing factors such as the initial neutrino-nucleon NC interaction, the nuclear model, the final-state interaction model, the nucleus deexcitation, and the secondary interaction on final-state particles.
@@ 1 Introduction
The neutral-current (NC) interaction between atmospheric neutrinos () and nuclei is a prominent channel in large water-Cherenkov (wCh) and liquid-scintillator (LS) detectors, particularly for neutrinos in (0.1 20) GeV range [2]. These NC interactions play a crucial role in contributing to irreducible backgrounds in large neutrino detectors, encompassing searches for rare events like the diffuse supernova neutrino background (DSNB) [3, 4, 5, 6, 7], proton decay [8, 9, 10, 11], and neutrinos form the dark matter annihilation [5, 12, 13]. Meanwhile, their observable energy range can extend into the reactor neutrino observation window, mimicking the reactor neutrino signal and further enhancing irreducible backgrounds.
@@
Reactor neutrinos play a crucial role in the field of neutrino physics, especially in the measurement of neutrino oscillations [14, 15, 16]. Future large liquid-scintillator (LS) detectors, exemplified by the Jiangmen Underground Neutrino Observatory (JUNO) [17], exhibit significant potential to attain sub-percent precision in measuring neutrino oscillation parameters [18] and determining neutrino mass ordering [17]. This potential stems from their capability to detect reactor inverse-beta-decay (IBD) reactions, , characterized by a prompt positron signal and a delayed neutron capture signal. LS detectors, featuring excellent neutron tagging, inherently offer high efficiency in IBD event selection. Nonetheless, for achieving highly precise measurements of reactor , careful consideration of -C NC interactions is essential.
@@
In the preceding paper, denoted as “the preceding” paper [1], we systematically explored the NC interaction of -C in LS detectors. For our calculations, we utilized the latest atmospheric neutrino fluxes at the JUNO site from the Honda group [19, 20]. To determine the rates and the spectra for both inclusive and exclusive channels of -C NC interactions within the LS, we adopted a two-step calculation approach. Initially, we employed the advanced neutrino interaction generators GENIE [21] and NuWro [22] to calculate neutrino-C interactions. Subsequently, the TALYS [23] package was used to predict the deexcitation of the residual nucleus. The preceding paper systematically explored the NC background concerning DSNB and proton decay. Building on this foundation, the present work continues the systematic investigation of NC background relevant to experimental research on the IBD signals of . This investigation focuses on the visible energy range below 100 MeV, specifically targeting reactor neutrinos and DSNB.
In contrast to the preceding paper, this work introduces significant updates. Firstly, it integrates advancements by implementing five sophisticated, data-driven neutrino generator models from GENIE (3.0.6) and NuWro (19.02). Secondly, these new models are used to explore the characteristics of -C interaction, including the initial -nucleon interaction within C and final-state information to demonstrate the model variations. This investigation elucidates the impact of the initial nuclear state, final-state interaction (FSI) model, and deexcitation process on these characteristics. Thirdly, GEANT4 [24] is employed to simulate the propagation of final-state particles, converting their kinetic energies into visible energies in the LS. The work estimates the impact of secondary interactions of final-state particles on neutron multiplicity and the exclusive channels associated with the NC background. Finally, this comprehensive approach allows for obtaining reliable predictions of the NC background in the observation of reactor neutrinos, DSNB, as well as neutrinos from the dark matter annihilation.
@@
The structure of the remaining sections in this paper is outlined as follows. In Sec. 2, we present a comprehensive overview of the calculation methodology, discussing the evolution of neutrino generator models from early to recent times, the TALYS-based deexcitation model, and the GEANT4-based detector simulation. Sec. 3 focuses on highlighting key characteristics derived from the simulation of the -C interaction, including discussions on the initial -nucleon interaction and final-state information. Moving to Sec. 4, we present the final results of the NC background. Finally, in Sec. 5, we provide a summary of our main results and conclude the paper.
2 Methodology for -C NC Interaction Prediction
Fig. 1 illustrates the potential NC interaction of with C in the LS detector. It includes the initial -nucleon NC interactions and nuclear effects, such as nuclear structure, many-body effects, FSI, and deexcitation of the residual nucleus. Therefore, compared to the NC interactions between and free nucleons, the calculation of -C NC interaction processes should include the following aspects. First, since the individual nucleon participating in the NC interaction is confined in the nucleus, the nuclear structure should be taken into account. In Fig. 1, the simple shell model of C [25, 26, 27] is used as an example to demonstrate the nuclear structure. Second, as the energy of the incident neutrino increases from 100 MeV to 10 GeV or even higher, the dominant contribution to the cross-section generally comes from interaction processes, including quasielastic scattering (QE), two-particle two-hole interaction (), coherent and diffractive production (COH), nuclear resonance production (RES) and deep inelastic scattering (DIS). Third, a more comprehensive study must also incorporate FSI and the deexcitation of excited residual nuclei, which involve the ejection of nucleons, ’s, and potentially even heavy particles.
2.1 Neutrino Generator Models: From Early to Recent Times
In the preceding paper [1], we have employed six representative neutrino generator models to calculate the neutrino-nucleus interactions using GENIE (2.12.0) and NuWro (17.10) generators. These models include one from GENIE, referred to as Model-G, and five from NuWro, denoted as Model-N for . Tab. 1 summarizes the main features of these six typical models at the early time stage. Detailed information and descriptions of these models can be found in the preceding paper [1], with a summarized overview provided below.
To investigate the impact of the axial mass , which is used to parameterize the nucleon axial-vector form factor in the QE process and is subject to uncertainty, different models make distinct choices for . Model-G adopts a default setting of GeV [28]. In NuWro, three different values are utilized: GeV for Model-N for , 1.03 GeV [29] for Model-N2, and 1.35 GeV [30] for Model-N3.
Concerning nuclear models, Model-G integrates the relativistic Fermi gas model with “Bodek-Ritchie” modifications (BRRFG) [31]. On the other hand, NuWro provides various options for describing the target nucleus, including both the local Fermi gas (LFG) and the spectral function (SF) approach, used in Model-N for and Model-N5, respectively. It’s worth noting that there is a typo in Ref. [1, 32]: Model-N for employs the LFG model instead of the RFG model.
To account for the two-body current effects, Model-N4 includes the transverse enhancement (TEM) model [33] of the meson exchange current from NuWro.
The method used to predict -C NC interaction in this work is inherited from the preceding paper [1]. As our understanding of neutrino-nucleus interaction models and neutrino generators advances, these six models used in the preceding paper are reexamined in the following.
The long-standing GENIE project has reached a significant milestone with the release of GENIE v3 [37], which includes substantial improvements in neutrino interaction models, compared to GENIE v2. Therefore, the GENIE generator has been updated from 2.12.0 to 3.0.6 in our calculations to predict -C interactions. Through a comparison of predictions of -C charge current (CC) interactions using the Transverse Kinematic Imbalance (TKI) [38, 39] by GENIE v3 generator with MINERA data [40], the prediction from a combination setup involving LFG and the INTRANUKE hN 2018 model (hN2018) exhibits greater consistency with the MINERA CC data [41].
The NuWro generator has been updated from (17.10) to (19.02) in our calculations, and the cross-section configurations from NuWro (19.02) have been validated by the TKI with MINERA data [40]. The TKI results [41] indicate that the predictions from the models with nuclear model LFG or SF are more consistent with the data.
It is increasingly evident that the large value of 1.35 GeV for QE is derived without accounting for the contribution from the two-body current interactions [44, 45, 33]. However, a direct measurement of the interaction has not yet been realized, leading to a significant model-dependent uncertainty. For instance, the TEM model implemented in NuWro can simulate the NC interactions for neutrinos but is absent in antineutrino processes. Additionally, in NuWro, there are no implements for NC interactions using the Nieves model [44].
@@ 2.1 Neutrino Generator Models: From Early to Recent Times
| Model-G1 (G) | GENIE (2.12.0) | 0.99 | BRRFG | hA | |
@@
| Model-G2 | GENIE (3.0.6) | 0.96 | LFG | ✓(EP) | hN2018 |
| Model-G3 | GENIE (3.0.6) | 0.96 | LFG | ✓(EP) | hA2018 |
| Model-G4 | GENIE (3.0.6) | 0.96 | BRRFG | ✓(EP) | hN2018 |
@@
In Tab. 1, we present a comparison of key features between early and recent models incorporated into our calculations to demonstrate the variations in neutrino interactions. In our present work, we have introduced five typical models, including three new models (i.e., Model-G2, Model-N6, and Model-N7) in response to constraints from external data, as mentioned earlier. Model-G2 incorporates the G18_10b_02_11a configuration from GENIE (3.0.6), combining the LFG framework for nuclear models and the hN2018 model for FSIs. Model-N6 and Model-N7 have configurations similar to Model-N1 and Model-N5, respectively.
In addition to Model-G2, we have expanded Model-G3 and Model-G4 from GENIE (3.0.6) to explore the effects of different FSI and nuclear models on the predictions of NC events. Model-G3 uses the INTRANUKE hA 2018 model (hA2018) to assess the impact of FSI uncertainty, while Model-G4, employing the BRRFG model [31], investigates the uncertainty arising from different nuclear models. Parameters such as GeV for QE, the Empirical (EP) model [42] for interactions, and the Berger-Sehgal (BS) model [43] for resonance pion production remain the same for Model-G for . For clarity, we have renamed “Model-G” to “Model-G1.”
Despite the variations introduced in the GENIE and NuWro models, it is important to note that the current model variations are limited. Firstly, recent models primarily focus on variations related to QE interactions and do not fully explore variations related to RES, COH, and DIS physics processes. Secondly, in addition to GENIE and NuWro used in our calculations, we plan to include other generators such as GiBUU [47] and NEUT [48].
@@ 2.2 TALYS-based Deexcitation Model of Residual Nucleus
The nuclear deexcitation involves additional rays, protons, neutrons, or other heavier projectiles. The production of neutrons and unstable residual nuclei during the deexcitation stage is crucial for tagging and reducing potential NC background in LS detectors. However, Monte Carlo neutrino generators typically lack information on the deexcitation of residual nuclei that may remain excited. The implementation of the deexcitation of the residual nucleus from NC interactions between atmospheric neutrinos and C follows the methodology outlined in the preceding paper [1] and must carried out separately from the neutrino generation.
The nuclear structure is specified using a simplified statistical configuration from the nuclear shell model of C [25, 26, 27], as depicted in Fig. 1. Additionally, we neglect the potential configuration of the C nucleus arising from nucleon pairing correlation between the and shells, given the relatively small energy gap of around 3-4 MeV[25, 26]. This effect will be included in our future work, incorporating more sophisticated shell model calculation. The excited states of residual nuclei are determined by considering the disappearance of one or more nucleons (either protons or neutrons) from the shell of the C ground state. Note that, in the simplified shell model configuration, if one or more nucleons disappear from the shell, these nucleons do not impact the state of the residual nucleus. The cases of one or two nucleons disappearing from C in shown in Table II of the preceding paper [1]. This table includes the corresponding probabilities of possible configurations together with the excitation energies of these residual nuclei. There are typos for the configuration probabilities for B and the corrected information is 4/9, 4/9 and 1/9 for MeV, respectively.
After determining the excited state of the residual nucleus with the excitation energy , TALYS (1.8) utilizes this information to simulate all potential deexcitation channels for the nucleus and obtain their deexcitation along with the energy spectra of the associated final-state particles. In our calculations, all residual nuclei with a mass number larger than five have been considered. For the dominant residual nucleus, C, B, and B, when they remain at 23 MeV, approximately 10%, 70%, and 20% of the deexcitation channels involve neutron production, respectively. The impact of deexcitation on NC background predictions will be discussed later.

### v2

# Neutral-current background induced by atmospheric neutrinos at large liquid-scintillator detectors: III. Comprehensive prediction for low energy neutrinos

@@ Abstract
Atmospheric neutrinos play a vital role in generating irreducible backgrounds in liquid-scintillator (LS) detectors via their neutral-current (NC) interactions with 12C nuclei. These interactions may affect a wide range of research areas from the MeV to GeV energy range, such as the reactor and geo neutrinos, diffuse supernova neutrino background (DSNB), dark matter, and nucleon decay searches. In this work, we extend our preceding paper [1], by conducting a first-time systematic exploration of NC backgrounds as low as the MeV region of reactor and geo neutrinos. We utilize up-to-date neutrino generator models from GENIE and NuWro, a TALYS-based nuclear deexcitation package and a GEANT4-based detector simulation toolkit for our complete calculation. Our primary focus is to predict the NC background for experimental searches of inverse-beta-decay signals below the 100 MeV visible energy. In order to have deeper understanding of the characteristics of atmospheric neutrino NC interactions in LS, we investigate the model dependence of NC background predictions by using various data-driven models, including the initial neutrino-nucleon interactions, nuclear ground-state structure, final-state interactions, nuclear deexcitation processes, and secondary interactions of final-state particles.
@@ 1 Introduction
The neutral-current (NC) interactions between atmospheric neutrinos () and nuclei are important signatures of the large water-Cherenkov (wCh) and liquid-scintillator (LS) detectors, especially for the neutrinos in the energy range of (0.1 20) GeV [2]. The products of NC interactions include the daughter nuclei, nucleons, mesons, -rays and invisible neutrinos. These final-state particles contribute to the signals of single, time-correlated double or triple-coincident events in either wCh or LS detectors. It offers an alternative way to explore the GeV neutrino interaction mechanism and enhance our understanding of these interactions [3, 4]. Meanwhile, the NC interactions significantly contribute to the irreducible backgrounds, encompassing the searches for rare events such as the diffuse supernova neutrino background (DSNB) [5, 6, 7, 8], proton decay [9, 10, 11, 12], and neutrinos from the dark matter annihilation [7, 13, 14]. In addition, the observable energy range of the NC interactions can overlap with the reactor neutrino energy observation window [15, 16, 17], mimicking the reactor neutrino signal and further escalating irreducible backgrounds. Future large liquid-scintillator (LS) detectors, exemplified by the Jiangmen Underground Neutrino Observatory (JUNO) [18], exhibit significant potential to attain sub-percent precision in measuring the neutrino oscillation parameters and determining neutrino mass ordering. This potential is from their capability to detect reactor inverse-beta-decay (IBD) reactions, , characterized by a prompt positron signal and a delayed neutron capture signal. LS detectors, featuring excellent neutron tagging, inherently offer high efficiency for the IBD event selection. Nonetheless, to achieve highly precise measurements of reactor , careful consideration of -12C NC interactions is essential.
@@
In our preceding paper [1], denoted as “the preceding” paper, we have systematically calculated the NC interactions of -12C in large LS detectors. For our calculation, we have utilized the fluxes at the JUNO site from the Honda group [19, 20], and determined rates and spectra for both inclusive and exclusive channels of -12C NC interactions within the LS. In the the preceding paper, we have adopted a two-step calculation approach. Firstly, we have employed advanced neutrino interaction generators GENIE [21] and NuWro [22] to calculate neutrino-12C interactions. Subsequently, the TALYS [23] package is used to predict the deexcitation of the residual nucleus. This prediction method for the NC background is extensively used in various physical topics within the JUNO collaboration, including reactor neutrinos to measure neutrino oscillation parameters [24] and identify neutrino mass ordering [25], the prospects for detecting the DSNB [26], the annihilation of MeV dark matter in the galactic halo [27] and invisible decay modes of neutrons [28].
Expanding on this basis, in the current study, we firstly integrate additional sophisticated neutrino generator models from the updated versions of GENIE (3.0.6) and NuWro (19.02). Moreover, by incorporating a GEANT4-based detector simulation [29] for the second interactions of particles propagating in the LS medium, we develop a three-fold strategy for predicting NC interactions. All models are employed to examine the characteristics of -12C interactions, encompassing both the initial -nucleon interaction within 12C and final-state distributions to demonstrate model variations. This investigation not only elucidates the influence of various models of initial nuclear states, final-state interactions (FSI), and the nuclear deexcitation process on the calculation of -12C interactions, but also incorporates the effect of secondary interactions of final-state neutrons on the final tagged neutron multiplicity. Finally, the preceding paper has delved into the implication of the NC background concerning the searches for the DSNB and proton decay. In contrast, this work extends to much lower energies to fit with the reactor and geo neutrino studies. Furthermore, with the introduction of new data-driven models, a discussion on the NC background for DSNB is included, providing a comparative analysis with the predictions from previous models.
@@
The structure of the remaining sections in this work is outlined as follows. In Sec. 2, we present a comprehensive overview of the calculation methodology, discussing the evolution of neutrino generator models from early to recent times, the TALYS-based deexcitation model, and the GEANT4-based detector simulation. Sec. 3 focuses on highlighting key characteristics derived from the simulation of the -12C interaction, including discussions on the initial -nucleon interaction and final-state information. Moving to Sec. 4, we present the final results of the NC background. Finally, in Sec. 5, we summarize our main results and conclude the paper.
2 Methodology for -12C NC Interaction Prediction
Our current comprehension of the interaction mechanism mainly relies on model predictions, especially for NC interactions between the and nuclei. Considering the primary energy range of is around the GeV scale, the NC interaction mechanism between the and nuclei is exceptionally complex, presenting numerous models each with its own set of challenges. Figure 1 illustrates the schematic procedure for calculating the NC interactions of with 12C. This includes the initial-state -nucleon NC interaction and various aspects of nuclear effects.
For the initial-state interaction, as the energy of the incident neutrino increases from 100 MeV to 10 GeV or even higher, the dominant contribution to the cross-section typically originates from different interaction processes, including quasi-elastic scattering (QE), coherent and diffractive production (COH), nuclear resonance production (RES) and deep inelastic scattering (DIS). These processes are explained by using parameterization of form factors, with the primary uncertainty lying in the axial form factors. For instance, in the current parameterized models of the NCQE interaction mechanism between the and nucleon, the axial mass constitutes one of the most challenging parameters for experimental determination. Despite numerous experiments measuring , it exhibits substantial measurement uncertainties. Present measurements estimate the value of in the range of around (1 1.4) GeV [30, 31, 32].
Moreover, in the -nucleus interactions, one should consider nuclear effects including the binding energy, nuclear ground state structure, many-body effects, FSI, deexcitation of the residual nucleus, and others. However, the nuclear effect models pose the following challenges:
The description of nuclear binding effects fluctuates across different models and requires additional exploration. The contemporary models, including Relativistic Fermi Gas (RFG), Local Fermi Gas (LFG), and Spectral Function (SF), are employed to characterize the nucleus ground state. However, these models may exhibit inconsistencies in describing the momentum distribution of nucleons bound within the nucleus.
The inclusion of two-body current effects in QE, referred to as two-particle two-hole interactions () or meson exchange currents (MEC), allows for the reproduction of MiniBooNE results with a smaller axial mass value, consistent with the world average obtained from other measurements [34, 35, 33]. However, a direct measurement of the interaction remains unattained, resulting in a considerable model-dependent uncertainty.
Nucleons involved in the neutrino initial-state interactions possess a high probability of interacting with surrounding nucleons within the nucleus, leading to FSI. However, discrepancies exist among various models in depicting this process. Moverover, post-FSI, the state of the residual nucleus and potential deexcitation processes remain incompletely addressed.
In addition to the primary interactions, predictions should also incorporate the secondary interactions, which occur when final-state particles interact within the detector after a primary interaction. These secondary interactions will influence the ultimate detector response and the particle identification in the detectors.
2.1 Neutrino Generator Models
In order to account for various contributions to the initial-state -nucleon interactions and manage the nuclear effects in neutrino-nucleus interactions, we typically use sophisticated Monte Carlo neutrino event generators. Besides GENIE and NuWro, which have been used in our calculations to demonstrate neutrino generator model dependence, other generators like GiBUU [36] and NEUT [37] are also available. These generators are carefully constructed and finely tuned according to available experimental data on interaction cross-sections. For example, predictions from GENIE and NuWro are evaluated against numerous experimental data like ANL, BNL, K2K, MiniBooNE and SciBooNE [21, 39, 38]. Nevertheless, as previously mentioned, the lack of sufficient experimental data constraints on both initial-state neutrino-nucleon interactions and nuclear effects results in predictions being subject to a considerable level of uncertainty. Consequently, neutrino event generators provide numerous possibilities for the physics parameters and nuclear effect models, offering a wide array of potential neutrino generator models for simulating neutrino interactions with the nuclei.
Given the existing focus on NC backgrounds within the visible energy range under 100 MeV, we adopt neutrino generator models that are predominantly centered on variations associated with the QE interactions between and 12C in LS. Specifically, eleven representative neutrino generator models are employed to evaluate the impact of different neutrino event generators, the axial mass in QE, nuclear ground state, effects, and FSI models on the predictions. Among these, six typical neutrino generator models are derived from GENIE (2.12.0) and NuWro (17.10). These models have been used in our preceding paper, with one from GENIE referred to as Model-G1 (G), and five from NuWro labeled as Model-N for .
GENIE and NuWro neutrino event generators persistently evolve their modern and universal event generator framework and tools to support neutrino experiments and provide novel comprehensive physics models for simulating neutrino interactions. For instance, GENIE has achieved a notable milestone with the release of GENIE v3 [40], which offers new physical tunes of neutrino interaction cross sections based on data from neutrino scattering experiments such as neutrino and antineutrino charge-current (CC) data on Carbon from MiniBooNE, T2K and MINERA [41, 42]. This signifies a shift from the prior approach in GENIE v2, which depends on a single ”Default” model with optional modeling elements. NuWro (19.02) significantly improves the nucleon cascade model [43], which is able to reproduce nuclear transparency data of electrons, particularly in the energy range crucial for neutrino-nucleus scattering physics. Hence, our calculation incorporates the updated GENIE (3.0.6) and NuWro (19.02) for predicting the interactions. We employ five new neutrino generator models from GENIE (3.0.6) and NuWro (19.02), including three from GENIE designated as Model-G for , and two from NuWro labeled as Model-N6 and Model-N7.
@@ 2.1 Neutrino Generator Models
| Model-G1 (G) | GENIE (2.12.0) | 0.99 | BRRFG | A | |
@@
| Model-G2 | GENIE (3.0.6) | 0.96 | LFG | ✓(EP) | N |
| Model-G3 | GENIE (3.0.6) | 0.96 | LFG | ✓(EP) | A |
| Model-G4 | GENIE (3.0.6) | 0.96 | BRRFG | ✓(EP) | N |
@@
In Tab. 1, we have compared the key features between these early adopted and recent models in our new calculation to illustrate the variations in neutrino interactions. The principal characteristics of these neutrino generator models and the reasoning behind their selection are summarized as follows.
- Initial neutrino-nucleon interaction:
In the initial stage, we scrutinize the impact of the value on the cross sections of QE interactions. Different models utilize unique choices for . For GENIE, Model-G1 employs a default setting of GeV, determined from the deuterium measurements [30]. Similarly, Model-G with adopts a very close value of GeV and these new models incorporate the G18_10a/b configurations and have been tuned by neutrino scattering data [42]. In NuWro, three different values are used: GeV for Model-N with , 1.03 GeV from the world average [31] for Model-N with , and 1.35 GeV from the MiniBooNE data [32] for Model-N3.
- Nuclear model:
In GENIE and NuWro, various nuclear models such as RFG, LFG, and SF are available to determine the momentum and binding energy of the target nucleon. To investigate the impact of the uncertainty from nuclear models on the prediction, Model-G1/4 integrates the RFG model with ”Bodek-Ritchie” modifications (BRRFG) [44], while Model-G2/3 uses LFG. Regarding the NuWro generator, LFG and SF are used in Model-N for and Model-N5/7, respectively.111It is important to note that there is a typo in Refs. [1, 45]: Model-N for employs the LFG instead of RFG.
- Inclusion of the effect:
Not all the models from GENIE and NuWro integrate models to contemplate the two-body current effects on NCQE. For our selection, Model-N4 features the transverse enhancement (TEM) model of the meson exchange current to predict the interactions, derived from fitting the electron scattering data [33]. In other NuWro models, a different implementation is used only for CC interactions, by using the Nieves model [34], a theory-based model commonly employed in the neutrino analyses. GENIE (3.0.6) models adopt the empirical (EP) multi-nucleon model with implementations of the corresponding Valencia models by Nieves et al. [46, 34].
- FSI model:
Describing the effect of FSI is crucial yet challenging for neutrino-nucleus interactions. In the early models of GENIE (2.12.0), FSI is simulated using the intranuclear cascade (INC) A model, an empirically tunable model primarily driven by hadron-nucleus scattering data [21]. Within GENIE (3.0.6), another FSI model, the INC N is available, which features a full cascade approach. To probe the impact of various FSI models on the predictions, Model-G1/3 employs the A model, whereas Model-G2/4 uses the N model for comparison. NuWro (17.10) models adopt the default FSI model [38], which is described in a framework of the INC model [47]. The specific dynamics is taken from the Oset model [48]. This NuWro FSI model is compared to the NC production data. NuWro (19.02) has upgraded the FSI model, in which the major improvements include an effective nucleon density considering nucleon-nucleon correlations, implementation of elastic nucleon-nucleon scattering angular coefficients, new inelastic nucleon scattering cross sections, single pion production fractions, and corrections to in-medium inelastic nucleon scattering cross sections [43].
In addition to the QE and interactions, all the other interaction processes such as RES, COH and DIS are also integrated into our calculations, in which the generator default settings are applied to all these processes. For instance, the Rein-Sehgal model [49, 50] is applied in Model-G1 to describe RES and COH, while a model using modifications suggested by the Bodek-Yang model [51] is utilized for DIS. The same cross section models of RES, COH and DIS are used in Model-G2/3/4, with the Berger-Sehgal model [52] for resonance pion production.
@@ 2.2 TALYS-based Deexcitation Model of Residual Nucleus
The nuclear deexcitation involves additional rays, protons, neutrons, or other heavier projectiles. The production of neutrons and unstable residual nuclei during the deexcitation stage is crucial for tagging and reducing potential NC background in LS detectors. However, Monte Carlo neutrino generators typically lack information on the deexcitation of residual nuclei that may remain excited. The implementation of the deexcitation of the residual nucleus from NC interactions between atmospheric neutrinos and 12C follows the methodology outlined in the preceding paper and must carried out separately from the neutrino generation.
The nuclear structure is specified using a simplified statistical configuration from the nuclear shell model of 12C [53, 54, 55], as depicted in Fig. 1. Additionally, we neglect the potential configuration of the 12C nucleus arising from nucleon pairing correlation between the and shells, given the relatively small energy gap of around 3-4 MeV[53, 54]. This effect will be included in our future work, incorporating more sophisticated shell model calculation. The excited states of residual nuclei are determined by considering the disappearance of one or more nucleons (either protons or neutrons) from the shell of the 12C ground state. Note that, in the simplified shell model configuration, if one or more nucleons disappear from the shell, these nucleons do not impact the state of the residual nucleus. The cases of one or two nucleons disappearing from 12C in shown in Table II of the preceding paper. This table includes the corresponding probabilities of possible configurations together with the excitation energies of these residual nuclei. There are typos for the configuration probabilities for 10B∗ and the corrected information is 4/9, 4/9 and 1/9 for MeV, respectively.
After determining the excited state of the residual nucleus with the excitation energy , TALYS (1.8) utilizes this information to simulate all potential deexcitation channels for the nucleus and obtain their deexcitation along with the energy spectra of the associated final-state particles. In our calculations, all residual nuclei with a mass number larger than five have been considered. For the dominant residual nucleus, 11C, 11B, and 10B, when they remain at 23 MeV, approximately 10%, 70%, and 20% of the deexcitation channels involve neutron production, respectively. The impact of deexcitation on NC background predictions will be discussed later.